The path effort of the following circuit is 26/9 Pro C خواب خط
Answer #1

False because first 3nand gates consider as one case logical path effort for this case is 4/3 since they are 2input NAND gate next we have not gates consider 3 NOT gates as one case logical path effort for this one next we have And gate it consider as one case it divide as NAND gate with not gate NAND have 3inputs so logical path effort for this is 5/3 so total logical path effort is= 4/3*1*5/3*1=20/9